OpenStudy (anonymous):

Select the set of equations that represents the following situation: Mike invested $199. for one year. he invested part of it at 5% and the rest at 7%. At the end of the year he earned $10.75 in interest. How much did Mike invest at each rate of interest? can you please help me

OpenStudy (anonymous):

can anyone help me

OpenStudy (kropot72):

Let p = part amount Remaining amount r = 199 - p Interest on p = p * 0.05 Interest on r = (199 - p) * 0.07 Total interest = (p * 0.05) + (199 - p) * 0.07 Do you follow so far?

OpenStudy (anonymous):

yes and the solution would be for both 5% and 7%

OpenStudy (anonymous):

you still there

OpenStudy (anonymous):

kropot can you help me finish this one

OpenStudy (anonymous):

thank you kropot, still waiting

OpenStudy (kropot72):

Total interest = 10.75 = 0.05p + (199 * 0.07) -0.07p 10.75 = -0.02p + 13.93 0.02p = -10.75 + 13.93 = 3.18 p = 3.18/0.02 = $159 The amount invested at 5% is $159 and the amount invested at 7% is 199 - 159 = $40

OpenStudy (anonymous):

I got this too, its just i did not come up with it the way that you did and i wanted to learn how to put this into an equation. Thank you Thank you!!!!!! soooo much.

OpenStudy (kropot72):

Good work!!! You're welcome :0